# RiskLabs: Predicting Financial Risk Using Large Language Model Based on Multi-Sources Data

### v1

@@ 2.3.1 Audio Encoding
To describe the Audio Encoding in more detail, we let the raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample. Each audio frame will be converted into a vector representation:
$e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i}).$ (1)
@@
Therefore, we obtain the audio embeddings $E_{ac}=\{e_{ac}^{1},e_{ac}^{2},\ldots,e_{ac}^{n}\}$. The shape of the $E_{ac}$ is 520 $\times$ 512, where 520 is the maximum number of audio files amongst all companies and 512 is the dimension of the resulting transform from the model for a single audio frame. Earnings conference calls with less than 520 audio frames ($n<520$) have been zero-padded for uniformity in input matrix size.
@@
$\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$ (2)
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (3)
@@
$\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$ (4)
@@
2.3.2 Transcript Encoding
The overall process is similar to Audio Encoding. We first use SimCSE Gao et al. (2021) to extract the vector representation of each sentence in earnings conference transcripts. SimCSE is a sentence-level representation as word/token-level representations are too difficult and document-level embedding is too sparse and loses a lot of information. SimCSE is a Siamese neural network architecture that learns to embed pairs of sentences into a shared space where similar sentences are mapped close together and dissimilar sentences are mapped far apart. We let raw transcripts as $T_{c}\ =\{t_{c}^{1},a_{c}^{2},\ldots,t_{c}^{n}\}$ where $t_{c}^{i}$ represents the $i^{th}$ sentence in the transcript. Therefore, each sentence will be converted into a vector representation:
$e_{tc}^{i}=\text{SimCSE}(t_{c}^{i}).$ (7)
@@ 2.3.2 Transcript Encoding
Same with (2)-(5), the MHA is applied to $E_{tc}$ to get $T_{tc}=\{t_{tc}^{1},t_{tc}^{2},\ldots,t_{tc}^{n}\}$ with dimension 520 $\times$ 768. Then, the average pooling layer is applied to $T_{tc}$:
$T_{t}=\text{AveragePooling}(T_{tc})$ (8)
@@
2.3.3 Earnings Conference Call Analyzer
An Earnings Conference Call is a teleconference or webcast in which a public company discusses its financial results for a reporting period, typically a quarter or fiscal year. During the call, senior company executives, including the CEO and CFO, present key financial figures, performance metrics, and strategies. They also provide insights into the company’s future outlook, including forecasts and potential challenges. Following the presentation, analysts and investors have the opportunity to ask questions.
In additional of the document structure, Earnings Conference Call within the same sector often exhibits similar discussion topics. For instance, in the energy sector, topics frequently revolve around oil and gas price fluctuations, regulatory changes, and environmental sustainability efforts. Earnings Conference Call in the consumer sector typically centers around consumer behavior trends, e-commerce growth, supply chain management, and product innovation.
To obtain insights from an Earnings Conference Call on how it might influence future market volatility, our approach encompasses two primary steps. (See Figure 2) Initially, we summarize the entire earnings conference call to capture a broad understanding of the company’s performance and future outlook. Following this, the second step involves a meticulous examination of specific points of interest to investors, for example focusing on detailed financial metrics and performance indicators mentioned in the call.
Figure 2: This figure visualizes the mechanism of the earnings conference call analyzer. It takes earnings conference calls as input. There is a two-step analysis: first, it summarizes the primary ideas. Second, it examines specific financial metrics and events mentioned in the call. At the top middle section, the illustration shows the analyzer segmenting the earnings conference call into smaller parts for more accurate summarization. The middle section demonstrates how the analyzer extracts and evaluates information from the earnings conference call. After this comprehensive analysis, the model produces a detailed analysis of the earnings conference call.
However, when attempting to summarize or retrieve information from earnings conference call transcripts using LLMs, directly inputting the entire document can pose significant challenges. The lengthy nature of these documents, often extending to thousands of words, along with the inclusion of sentences ranging from crucial insights to irrelevant remarks such as greetings, can affect the LLM’s ability to effectively comprehend and analyze the content of the earnings conference call.
Therefore, in the proposed Earnings Conference Call analyzer, it begins by segmenting the call transcripts based on the average length attributed to each topic. By organizing the text in this manner, we create distinct embeddings for paragraphs that revolve around similar ideas or themes. This strategy significantly enhances the model’s efficiency in information processing at later stages.
After dividing the lengthy earnings conference call transcript into several smaller sections, we use LLM to summarize each individual chunk. Since the content within each chunk is more homogeneous and concentrated around a single theme, the LLM can more effectively grasp the core idea. (See Figure 2 Green part) At the same time, to further combine the analyzed content with other extracted features, we use LLM to extract embedding from the analyzed earnings conference call as the analysis feature:
$T_{l}=\text{LLM}(\text{analysis of earnings conference call})$ (9)
After generating summaries for each segmented chunk of the earnings conference call, we proceed to aggregate these individual summaries and make a second summarization again. The initial summaries distill the essence of each chunk, stripping away extraneous details and focusing on the core information. When these condensed summaries are combined and summarized again, the reduced length facilitates an even more efficient and effective summarization by the LLM. Consequently, this two-tiered approach to summarization not only enhances the manageability of the text for the LLM but also ensures that the final summary is both comprehensive and succinct, capturing the most critical insights in a more accessible format.
In addition to summarization, the earnings conference call analyzer delves deeper into analyzing the transcript. We simulate the analytical process of financial analysts as they examine an earnings conference call. Experts will prioritize certain information, meticulously analyzing key financial metrics and performance indicators to gauge the company’s current state and future prospects.
To simulate this analysis process, we begin by identifying core topics analysts typically focus on in earnings conference call, such as dividends, earnings, and costs. Then we use LLM to locate and extract paragraphs mentioning these topics. However, accurately extracting this information from the dense narrative of earnings conference calls can pose a challenge. Because LLM searches are based on similarity, comparing a single-word topic to an entire document presents a scale issue, and a single word may lack the necessary context, thus diminishing the extraction capability. To enhance the precision of information extraction, we formulate a variety of questions related to each topic but phrased differently, ensuring a broad coverage and increasing the likelihood of accurately pinpointing relevant sections. For instance, regarding the dividends topic, we pose questions such as: (1) "Have there been any changes in the stock dividends, and if so, at what rate have the dividends increased or decreased?" (2) Does the company Board expect to increase the stock dividends in the next future? (3) How does the dividend yield compare to the peers?
After locating the segments where these topics are discussed, we extract the corresponding paragraphs. These paragraphs, however, may still contain irrelevant text such as preceding discussions. To address this, we use a contextual compression method from LLM, adept at distilling the paragraph down to its most relevant sentences, effectively obtaining the crucial data from the noise.
Our subsequent step involves determining how the company’s present position of each topic could influence future stock behavior. To this end, we consult financial experts to provide various potential scenarios within a specific topic and offer their professional insights on how each scenario could impact market responses. For example, regarding the dividend topics, the earnings conference call may discuss about increasing dividends and decreasing dividends in these two scenarios. Increasing dividends may drive positive returns and lower stock volatility. Conversely, a decrease in dividends might be perceived negatively by investors, leading to decreased returns and increased volatility in the stock’s price. Once again, we face the challenge of mapping sentences from earnings conference calls to various potential scenarios. Our approach involves a detailed, sentence-by-sentence comparison. (See Figure 2 Yellow part) We collect various expressions from multiple earnings conference calls, focusing on how speakers convey specific scenarios, like the increase of dividends. In this way, we can precisely determine the current scenario of the company, and then, guided by expert insights, we can gain an understanding of how the stock is likely to move in the coming days.
In summary, our analysis of the earnings conference call includes two parts: summarizing the documents to obtain an overview and extracting precise information for in-depth company analysis. Ultimately, we merge these two components to form an integrated earnings conference call encoder.
2.3.4 Additive Multi-modal Fusion
Given the model’s reliance on several inputs and diverse data types, we identify an effective fusion structure to integrate these features into the training process to ensure a balanced weighting among components. We use additive interactions to handle the representational fusion of different abstract representations. These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline Liang et al. (2022). Given the audio feature $T_{a}$, textual feature $T_{t}$ from the transcript, and $T_{l}$ from ECC analyzed text, additive fusion can be seen as learning a new joint representation:
$E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{l}+\epsilon$ (10)
where $w_{1}\in R^{512\times 512}$, $w_{2}\in R^{768\times 512}$ and $w_{3}\in R^{1024\times 512}$ are the weights learned for additive fusion, $w_{0}$ the bias term and $\epsilon$ the error term. $E$ is a vector with 512 as the final feature from the Earning Conference Call Encoder.
2.4 Time-Series Encoder
In this section, we mainly introduce how we can treat the time series data and how to capture the information from the time series data. Firstly, we collected daily VIX from CBOE. VIX reflects panic in the financial markets and is also an important indicator of market volatility. Furthermore, we downloaded the historical stock prices of each company in the S&P 500 and calculated the return and historical volatility for further analysis. We used two methods as our encoder to extract the information from the time series: Time-series embedding and the traditional econometrics model.
2.4.1 Time-Series Embedding
@@
For our time series input, we capture the VIX values from a 30-day period preceding the ECC release date. To distill meaningful features from this VIX series, we employ a Bidirectional Long Short-Term Memory (BiLSTM) network Siami-Namini et al. (2019):
$T_{v}=\text{BiLSTM}(\textbf{VIX})$ (11)
In this setup, the BiLSTM is configured with 64 hidden states. Consequently, the output $T_{v}$, representing the extracted time series features, has a dimensionality of 128, reflecting the bidirectional nature of the BiLSTM.
@@
2.4.2 Extract Relationship Among Multiple Response Variable Using Vector Auto-regression-based Method
@@
To achieve this, we attempt to find the relationship that links the two different stages: the $m$-day before and the $m$-day after. We will use the VAR(Vector Auto-Regression Model) to capture the linear relationship between the volatility of different terms. We set four different "$m$-day volatility"(3-day volatility, 7-day volatility, 15-day volatility, and 30-day volatility). We define the $\sigma_{-m,t}$ as the "$m$-day future volatility" at time $t-m$; the calculation method traces back $m-day$ from today and calculates the $m-day$’s standard deviation of returns.
@@
$P(\theta|Y,X)\propto P(Y|\theta,X)P(\theta|X)$ (13)
If we can determine the prior and the likelihood function, we can estimate the posterior distribution. We assume that the prior distribution $P(\theta)$ satisfies the Normal distribution based on the simple linear regression. Subsequently, we consider the Likelihood $P(Y|\theta,X)$. In the empirical study from Andersen et al. (2001), the unconditional distribution of the realized historical distribution of the volatility is highly right-skewed. Besides, Cizeau et al. (1997) tested the distribution historical volatility of S&P 500 stock from 1984 to 1996 and demonstrated that the pattern follows the log-normal distribution. Therefore, we took the log form of the original data. Given the preparation, we build the VAR model as follows:
$\left\{\begin{array}[]{c}log(\sigma_{3,t})=\alpha_{3}+\beta_{1,1}log(\sigma_{-3,t})+\beta_{1,2}log(\sigma_{-7,t})+\beta_{1,3}log(\sigma_{-15,t})+\beta_{1,4}log(\sigma_{-30,t})+u_{3,t}\\ log(\sigma_{7,t})=\alpha_{7}+\beta_{2,1}log(\sigma_{-3,t})+\beta_{2,2}log(\sigma_{-7,t})+\beta_{2,3}log(\sigma_{-15,t})+\beta_{2,4}log(\sigma_{-30,t})+u_{7,t}\\ log(\sigma_{15,t})=\alpha_{15}+\beta_{3,1}log(\sigma_{-3,t})+\beta_{3,2}log(\sigma_{-7,t})+\beta_{3,3}log(\sigma_{-15,t})+\beta_{3,4}log(\sigma_{-30,t})+u_{15,t}\\ log(\sigma_{30,t})=\alpha_{30}+\beta_{4,1}log(\sigma_{-3,t})+\beta_{4,2}log(\sigma_{-7,t})+\beta_{4,3}log(\sigma_{-15,t})+\beta_{4,4}log(\sigma_{-30,t})+u_{30,t}\end{array}\right.$
Here, $\sigma_{m,t}$ stands for the "$m$-day" future volatility at time $t$; $u_{m,t}$ is a White noise term, which $u_{m,t}\sim N(0,1)$; $\beta_{i,j}$ represents the linear relationship, which is the coefficient matrix of the VAR model, $\alpha_{m}$, which is the intercept term of the VAR model.
@@
Table 1: Bayesian-VAR results by applying the MCMC method
@@
In order to validate this methodology, we choose one of the stocks "TWTR," in our database as a case study, and then output the Bayesian regression results. Here, we used 250 days of data(from 2016-02-22 to 2017-02-15) as the model training period. Table 1 shows the training result. We observe that the $\hat{R}$ of all parameters is equal to 1, and the $ess\_tail$ value and $ess\_bulk$ are all over 7000. Typically, there is no specific argument about the number of ESS. Kruschke (2014) in the book mentioned that we need more ESS samples for the interested parameters. In the meantime, we also need to balance the number of ESS data and model training costs. Vehtari et al. (2021) in the paper recommends having the ESS number over 400, which should be dependent on practical experience and model, to get a stable Monte Carlo standard error.
Table 2: Statistics Results about AEP(Absolute Error Percentage) of the Experiment
@@
2.5 News-Market-Reactions Encoder

### v2

@@ 2.1.1. Audio Encoding
Audio embeddings are extracted using Wav2vec2 (Baevski et al., 2020) and then processed through a Multi-Head Self-Attention (MHSA) layer with pooling to obtain audio features $T_{a}$. To describe the Audio Encoding in more detail, we let the raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample. Each audio frame will be converted into a vector representation:
(1) $e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i}).$
@@
We obtain the audio embeddings $E_{ac}=\{e_{ac}^{1},e_{ac}^{2},\ldots,e_{ac}^{n}\}$. The shape of the $E_{ac}$ is 520 $\times$ 512, where 520 is the maximum number of audio files amongst all companies and 512 is the dimension of the resulting transform from the model for a single audio frame. Earnings conference calls with less than 520 audio frames ($n<520$) have been zero-padded for uniformity in input matrix size.
@@
(2) $\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$
(3) $\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$
@@
(4) $\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$
@@
2.1.2. Transcript Encoding.
Similar to Audio Encoding, this component first uses SimCSE (Gao et al., 2021) to generate vector representations for each sentence in the earnings call transcripts. These embeddings are then passed through an MHSA layer with pooling to capture text features $T_{t}$; We let raw transcripts as $T_{c}\ =\{t_{c}^{1},a_{c}^{2},\ldots,t_{c}^{n}\}$ where $t_{c}^{i}$ represents the $i^{th}$ sentence in the transcript. Therefore, each sentence will be converted into a vector representation:
(7) $e_{tc}^{i}=\text{SimCSE}(t_{c}^{i}).$
@@
Same with (2)-(5), the MHSA is applied to $E_{tc}$ to get $T_{tc}=\{t_{tc}^{1},t_{tc}^{2},\ldots,t_{tc}^{n}\}$ with dimension 520 $\times$ 768. Then, the average pooling layer is applied to $T_{tc}$:
(8) $T_{t}=\text{AveragePooling}(T_{tc})$
@@
2.1.3. Earnings Conference Call Analyzer.
This component utilizes an LLM to summarize the earnings call, extracting essential and relevant information and converting it into a text feature. We show this process in Figure 2.
In order to effectively summarize key information in lengthy ECC records, we employ a hierarchical summarization strategy. First, the entire document is divided into chunks, and then we use LLM to summarize each chunk individually. These individual summaries are then further summarized through LLM, resulting in a comprehensive summarization paragraph of the entire document. This two-layer approach ensures that both detailed and aggregate information is captured. In further, we use the OpenAI ‘text-embedding-3-small’ text embedding model to generate embeddings $T_{s}$ with size 1024 for summarized paragraph:
(9) $T_{s}=\text{Embedding}(\text{Concatenated}[chunk\ summaries;summary])$
Next, we aim to extract the most critical sentences containing relevant information from the entire transcript. To achieve this, we first convert the processed chunks into vector representations using the embedding model. Then, with input from financial experts, we design a list of questions about the ECC, which are stored in a ’Question Bank’ as a query list (see Appendix  LABEL:ssec:question). For each question in the ’Question Bank’, we retrieve the relevant chunks from the vectors individually. Next, we use the context compressor, in combination with the LLM chat, to quiz each extracted chunk and query on its relevance. This process filters out only the sentences useful for answering the questions. Finally, we synthesize the query-selected sentences into a coherent prompt, to which the large language model will generate responses as the final answers to the queries. These generated answers are considered as the extracted important sentences. We then concatenate all the generated sentences and feed them into a text embedding model to generate the text embedding $T_{f}$ with size 1024:
(10) $T_{f}=\text{Embedding}(\text{Concatenated}[selected\ sentences;])$
Table 1. Performance results on our proposed framework RiskLabs from different baseline models.
Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$ Multi-Task
Classical Method 0.713 1.710 0.526 0.330 0.284 / $\otimes$
LSTM 0.746 1.970 0.459 0.320 0.235 / $\otimes$
MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233 / $\otimes$
HAN 0.598 1.426 0.461 0.308 0.198 / $\otimes$
MRDM 0.577 1.371 0.420 0.300 0.217 / $\otimes$
HTML 0.401 0.845 0.349 0.251 0.158 / $\checkmark$
GPT-3.5-Turbo 2.198 2.152 1.793 2.514 2.332 0.371 $\checkmark$
RiskLabs 0.324 0.585 0.317 0.233 0.171 0.049 $\checkmark$
2.2. Time-Series Encoder
@@
For our time series input, we capture the VIX values from a 30-day period preceding the ECC release date. To distill meaningful features from this VIX series, we employ a Bidirectional Long Short-Term Memory (BiLSTM) network (Siami-Namini et al., 2019): $T_{v}=\text{BiLSTM}(\textbf{VIX})$. In this setup, the BiLSTM is configured with 64 hidden states. Consequently, the output $T_{v}$, representing the extracted time series features, has a dimensionality of 128, reflecting the bidirectional nature of the BiLSTM.
2.3. Relevant News Encoder
News significantly influences stock movements and serves as a valuable indicator of market trends. It encompasses a wide range of information, including macroeconomic indicators, industry trends, and company-specific events such as earnings reports, mergers and acquisitions, regulatory changes, and management shifts. For each stock, we compile relevant news into a single text string, which is then processed by an LLM to generate a news feature $T_{n}$.
2.4. Multimodal Fusion and Multi-Task Learning
The additive fusion method can be seen as learning a new joint representation:
(11) $E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{f}++T_{v}+T_{n}+\epsilon$
We concurrently model volatility prediction and VaR prediction using a multi-task framework. The multi-task prediction module is comprised of two separate single-layer feedforward networks, each responsible for predicting volatility (vol) and Value at Risk (var) values individually.
RiskLabs is a multi-task training model that trains and predicts the volatility of different terms and VaR. The individual stock volatility can be expressed as the natural log of the standard deviation of return prices $r$ in a window of ${\tau}$ days. We calculate the 3, 7, 15, and 30 trading days volatility using the equation:
(12) $v_{[d-\tau,d]}=\ln\left(\left(\frac{\sum_{i=0}^{\tau}(r_{d-i}-\bar{r})^{2}}{\tau}\right)^{\frac{1}{2}}\right)$
The notations used across the paper are discussed herein. Let $s\in S$ denote a stock, $c\in C$ be an earnings call for stock s. For each stock, there exist multiple earnings calls c that are held periodically. Each call c can be segmented into a set of $a_{c}^{i}\in A_{c}$ audio clips, and corresponding $t_{c}^{i}\in T_{c}$ text sentences for $i\in[1,N]$, where $N$ is the maximum number of audio clips in a call. For each stock, there exists a daily return denoted by $r_{i}=\ \frac{p_{i}-p_{i-1}}{p_{i-1}}$, where $p_{i}$ is the adjusted close price at the end of the day and r̄ is the average return over a period of ${\tau}$ days. We aim to develop a predictive regression function $f(c)\rightarrow v_{[d-\tau,d]}$
$v=F^{-1}(p)$
The $F(\cdot)$ is the cumulative loss distribution, $p$ is the percentile we set, and $v$ is the VaR. From the idea of Quantile Regression, we can have:
We train RiskLabs by optimizing multitask loss:
(13) $\mathcal{L}=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})+(1-\mu)\text{max}(q\times(v-\hat{v}),(1-q)(\hat{v}-v)).$
3. Experiment and Discussion
3.1. Experiment Setup
We compare our approach to volatility prediction to several important baselines as described in the Appendix A.2. We use GPT-4 for the Earnings Conference Call Analyzer and the News-Market Reactions Encoder. Throughout this process, we set the temperature parameter to 0. This ensures that the Large Language Models (LLMs) produce the most predictable responses, which aids in maintaining consistency in our experiments. For the overall training of the framework, we developed the code using PyTorch. Each Multi-Head Attention layer in the network comprises 8 individual heads, and the training process utilized batch sizes $b\in\{2,4,8\}$. We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$. The best hyper-parameters were kept consistent across all experiments, with the exception of the trade-off parameter $\mu$ which varied between the two tasks. We list additional details in Appendix A.1.
3.2. Performance Comparison (RQ1 & RQ2)
Table 2. Ablation Study: performance results of different modules.
Module $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$
Audio + Text 0.373 0.645 0.362 0.280 0.204 0.131
Audio + Text + Analysis 0.357 0.627 0.335 0.267 0.199 0.057
Audio + Text + Analysis + VIX 0.324 0.585 0.317 0.233 0.171 0.049
Table 3. Comparison of Value at Risk Predictions: AI Techniques vs. Traditional Financial Methods. The predefined VaR value is 0.05.
Method Prediction of VaR
Historical Method 0.016
Fully Connected Neural Network 0.044
LSTM 0.056
RiskLabs 0.049
The predefined VaR value is 0.05, meaning that the closer the model’s predictions are to 0.05, the better its performance. The result above shows that the prediction of VaR via applying the historical method is 0.016, which is significantly below the pre-defined percentile(5%). It indicates that the historical method overestimated the 95% VaR benchmark. Tracing back to 2016, we know there was a global financial crisis in 2015, and its effect lasted till the beginning of 2016. On January 20, 2016, the price of crude oil fell below \$27 a barrel; the DJIA index took a roller coaster from down 565 points to down 249 intraday. In February, the YTD (yield to Date) return came to -10.5%. These events, together with the sequelae of the 2015 stock market crisis, define 2016 as a risky year. Comparing 2016, 2017 will be much better. In January 2017, DJIA achieved a new historical height, landing above 20,000. The stock market experienced a boost with a 25% growth rate for DJIA, 19% for S&P 500, and 28% for Nasdaq. The market had strong confidence, and the VIX index in 2017 came to its historical lowest point. That explains the reason why the historical method may overestimate the 95% VaR benchmark, due to this method duplicating the extreme scenarios from 2016 to 2017, which leads to the extra estimation of financial risks. We show more results of the analysis and visualization in Appendix B.1.
3.3. Comparison Across Modules (RQ3)
In addition to overall performance, we evaluate the impact of different data combinations on predictive performance to better understand the relative contributions of data sources and each sub-module. We designed the ablation study as follows:
As shown in Table 2, the Audio + Text’ combination outperforms the baseline HTML model for 3-day forecasts. For longer periods—7, 15, and 30 days—RiskLabs achieves predictions closely aligned with HTML while still surpassing other baselines. Notably, with just Audio + Text’, our model is more streamlined than HTML, benefiting from a pre-trained model with extensive parameters that effectively map text and audio to vector representations. These vectors are processed through a multi-head self-attention mechanism for feature extraction, demonstrating the value of large-scale models in enhancing AI-driven risk prediction. Further improvements were observed upon integrating earnings conference call analysis and time-series data, leading to incremental gains in RiskLabs’ performance, especially for medium- to long-term forecasts (7, 15, and 30 days). These results suggest that earnings conference calls, as essential references for investors, have a more substantial impact on short-term risk volatility, while diverse information sources are crucial for improving accuracy in long-term forecasts. These findings underscore the contributions of each component within the RiskLabs framework, culminating in a robust predictive model.
3.4. Challenge and Opportunity
Our experimental findings demonstrate that leveraging large language models to integrate diverse information sources can significantly enhance the capability of our framework in predicting financial risks. With this in mind, we’ve expanded our data sources to include variables that could influence market volatility, such as news disseminated through social media. As detailed in Section 2.4, we utilize LLM to collect and analyze daily financial news. Initial small-scale experiments indicate that this addition can further improve the model’s performance. However, scaling up has introduced challenges: 1) the variable quality of news sources, with some containing misinformation, potentially introduces noise into our news database and affects audio model predictions; 2) there is a need to test the module’s effectiveness across a broader data spectrum. To address these challenges, we are implementing more nuanced steps to filter out low-quality news. Additionally, we are amassing newer data samples to enlarge our dataset, allowing for more comprehensive validation of our model’s efficacy. Furthermore, we will implement: a) Bayes-VaR method provides a probabilistic framework that allows for the incorporation of prior knowledge and the updating of beliefs as new evidence is presented. b) News-Market-Reactions Encoder c) ”Dynamic Moving Time Window” and d) a ”Time Decay Hyper-parameter” in RiskLabs. These features will enable more flexible training and forecasting on a daily basis, thus offering investors timelier and more accurate risk assessments. We described ongoing tasks as follows.
@@ 3.4. Challenge and Opportunity
3.4.1. Extract Relationship Among Multiple Response Variable Using Vector Auto-regression-based Method
@@
To achieve this, we attempt to find the relationship that links the two different stages: the $m$-day before and the $m$-day after. We will use the VAR(Vector Auto-Regression Model) to capture the linear relationship between the volatility of different terms. We set four different ”$m$-day volatility”(3-day volatility, 7-day volatility, 15-day volatility, and 30-day volatility). We define the $\sigma_{-m,t}$ as the ”$m$-day future volatility” at time $t-m$; the calculation method traces back $m-day$ from today and calculates the $m-day$’s standard deviation of returns.
@@
(15) $P(\theta|Y,X)\propto P(Y|\theta,X)P(\theta|X)$
If we can determine the prior and the likelihood function, we can estimate the posterior distribution. We assume that the prior distribution $P(\theta)$ satisfies the Normal distribution based on the simple linear regression. Subsequently, we consider the Likelihood $P(Y|\theta,X)$. In the empirical study from (Andersen et al., 2001), the unconditional distribution of the realized historical distribution of the volatility is highly right-skewed. Besides, (Cizeau et al., 1997) tested the distribution historical volatility of S&P 500 stock from 1984 to 1996 and demonstrated that the pattern follows the log-normal distribution. Therefore, we took the log form of the original data. Given the preparation, we build the VAR model as follows:
$\left\{\begin{array}[]{l}log(\sigma_{3,t})=\alpha_{3}+\beta_{1,1}log(\sigma_{-3,t})+\beta_{1,2}log(\sigma_{-7,t})+\beta_{1,3}log(\sigma_{-15,t})\\ +\beta_{1,4}log(\sigma_{-30,t})+u_{3,t}\\ log(\sigma_{7,t})=\alpha_{7}+\beta_{2,1}log(\sigma_{-3,t})+\beta_{2,2}log(\sigma_{-7,t})\\ +\beta_{2,3}log(\sigma_{-15,t})+\beta_{2,4}log(\sigma_{-30,t})+u_{7,t}\\ log(\sigma_{15,t})=\alpha_{15}+\beta_{3,1}log(\sigma_{-3,t})+\beta_{3,2}log(\sigma_{-7,t})\\ +\beta_{3,3}log(\sigma_{-15,t})+\beta_{3,4}log(\sigma_{-30,t})+u_{15,t}\\ log(\sigma_{30,t})=\alpha_{30}+\beta_{4,1}log(\sigma_{-3,t})+\beta_{4,2}log(\sigma_{-7,t})\\ +\beta_{4,3}log(\sigma_{-15,t})+\beta_{4,4}log(\sigma_{-30,t})+u_{30,t}\end{array}\right.$
Here, $\sigma_{m,t}$ stands for the ”$m$-day” future volatility at time $t$; $u_{m,t}$ is a White noise term, which $u_{m,t}\sim N(0,1)$; $\beta_{i,j}$ represents the linear relationship, which is the coefficient matrix of the VAR model, $\alpha_{m}$, which is the intercept term of the VAR model.
@@
Table 4. Bayesian-VAR results by applying the MCMC method
@@
In order to validate this methodology, we choose one of the stocks ”TWTR,” in our database as a case study, and then output the Bayesian regression results. Here, we used 250 days of data(from 2016-02-22 to 2017-02-15) as the model training period. Table 4 shows the training result. We observe that the $\hat{R}$ of all parameters is equal to 1, and the $ess\_tail$ value and $ess\_bulk$ are all over 7000. Typically, there is no specific argument about the number of ESS. (Kruschke, 2014) in the book mentioned that we need more ESS samples for the interested parameters. In the meantime, we also need to balance the number of ESS data and model training costs. (Vehtari et al., 2021) in the paper recommends having the ESS number over 400, which should be dependent on practical experience and model, to get a stable Monte Carlo standard error.
Table 5. Statistics Results about AEP(Absolute Error Percentage) of the Experiment
@@
3.4.2. News-Market-Reactions Encoder